Assign oxidation numbers to all elements in the following compound and ion:
a) NBr3

N__

Br__

b) MnO4-

Mn___

O___
Answer #1

Ans:- Oxidation number: - NBT NBr. X+ (-1)3 =0 I= +3 3 +3X=0 So-> N +3 Br - 1 (6) Mnou x +4+2) 4 = -1 De = -1 +8 x = +7 Mnoy